Use the Cauchy-Schwarz inequality to solve x^3-25x^2-4x+100=0, if we know
that all its roots are rational.
1
Expert's answer
2020-06-04T11:32:09-0400

"x^3 -25x^2- 4*x + 100 = 0"

imagine the left side as factors

"(x-2)(x-25)(x+2)=0"

each bracket may be equal 0. thats we have next three equations:

"x-2=0\\
x-25=0\\
x+2=0"

This equations have three roots: "x=2", "x=25", "x=-2"

This equation cannot be solved using the Cauchy-Schwartz inequality, since we have equality, and in the Cauchy-Schwarz inequality two variables are used, and we have one.